in a metal alloy, the ratio of zinc to copper is 3 to 15. If there are 630 pounds of copper, how much zinc is there?
Answer by stanbon(75887)

in a metal alloy, the ratio of zinc to copper is 3 to 15. If there are 630 pounds of copper, how much zinc is there?
------------------------
3:15 is the same as 3x:15x where 3x is zinc and 15x is copper.
----
Equation:
3x + 15x = 630
18x = 630
x = 35
---
Amt of zinc = 3x = 3*35 = 105 pounds
=========================================
Cheers,
Stan H.
==============